OpenStudy (anonymous):

2 cos^2o-1=0 Answer anyone??

OpenStudy (anonymous):

Is your equation \[2 cos^2 \Theta - 1 = 0\]

OpenStudy (anonymous):

If that is your equation, then this is a simple double-angle trigonometric identity. \[2cos^2 \Theta - 1 = cos(2 \Theta)\]

OpenStudy (tylerd):

so then cos(2theta)=0 ??

OpenStudy (anonymous):

are you solving for the angle?

OpenStudy (anonymous):

it can be solve directly actually...\[2\cos^2\theta-1=0\]\[2\cos^2\theta=1\]\[\cos^2\theta=\frac{1}{2}\]\[\cos\theta=\pm\sqrt{\frac{1}{2}}=\pm\frac{\sqrt{2}}{2}\]\[\theta=\cos^{-1}\left(\pm\frac{\sqrt{2}}{2}\right)\]\[\theta=\frac{k\pi}{4},~k=1,~3,~5,...~odd~integers\]

OpenStudy (anonymous):

cos ^2 theta=1/2